# Embedding Similarity Guided License Plate Super Resolution

### v2

@@ 5 Conclusion
[…]
The analysis of distortion maps underscores the model’s capability to mitigate blurring and compression artifacts while maintaining sharpness, a critical factor in generating high-quality SR images. By enforcing embedding distance minimization through the proposed PECL, the SR model achieves superior alignment between SR and HR license plates. This results in semantic consistency and improved recognition accuracy. Additionally, the loss function dynamically optimizes the contributions of pixel-wise and embedding-based losses during training, ensuring stability, adaptability, and robustness against varying degradation levels.
In summary, this work contributes to advancing LPSR by combining an innovative loss function with a robust architectural design. The proposed model demonstrates competitive performance, delivering visually accurate results compared to state-of-the-art methods, such as SwinIR and TBSRN, while enhancing the text recognition task.

### v3

@@ 5 Conclusion
This article presents a novel framework for license plate super-resolution (LPSR) that substantially outperforms existing methods by integrating a pixel and embedding consistency loss (PECL). The proposed approach, which leverages a Siamese network with contrastive loss, effectively balances pixel-level fidelity with embedding-level similarity, leading to consistent improvements in PSNR, SSIM, LPIPS, and OCR metrics. Experimental results demonstrate that our method not only recovers fine details and reduces blurring and compression artifacts but also enhances semantic consistency, as evidenced by superior text recognition performance on the CCPD dataset, achieving the highest exact match accuracy and the lowest error rates among competing methods.
Furthermore, qualitative comparisons across datasets reveal that while methods such as ESRGAN perform relatively well on the smaller-scale PKU dataset, our model consistently delivers more natural textures and sharper edges, thereby ensuring robust performance under diverse degradation conditions. The dynamic optimization of pixel-wise and embedding-based losses during training ensures stability and adaptability, ultimately yielding an SR output that is both visually accurate and task-aware. In summary, this work advances LPSR by combining an innovative loss function with a robust architectural design, demonstrating significant improvements over state-of-the-art techniques.